OpenStudy (anonymous):

Romero has the following scores on his math tests: 79, 86, 95, 88. How much would his mean change if the lowest test score is dropped? Round to the nearest whole number. a. 6 points b.5 points c. 4 points d.3 points

OpenStudy (anonymous):

d 3 points

OpenStudy (anonymous):

JHMann tke back you words

OpenStudy (anonymous):

(79+86+95+88)/4 = 87 and (86+95+88)/3 = 89.667 which rounds to 90 and 90-87 = 3

OpenStudy (anonymous):

I agree with JHMann...